# Heterogeneous Treatment Effects and

### v1

@@ 3 Framework
We introduce a framework that we use to analyze the relationship between HTEs—as estimated by treatment-covariate interactions—and the detection of causal mechanisms. Our framework is built upon the potential outcomes framework or Neyman-Rubin causal model (Neyman, 1923; Rubin, 1974). We denote treatment by $Z$. In order to consider HTEs with respect to pre-treatment moderators, denote the vector of measured pre-treatment covariates by $X=(X_{1},X_{2},...,X_{k})\in\mathbb{R}^{k}$. Clearly, it need not be the case that all $n$ covariates moderate the treatment effect.
A valid mediator, or mechanism representation, should: (1) be affected by treatment, $Z$, and (2) have a non-zero effect on the outcome. Importantly, a causal mediator could be affected by some covariate(s) $X_{k}$. This “arrow” from some $X_{k}$ to the value of the mediator is essential for using heterogeneous treatment effects to detect causal mechanisms. In the above example, the mediator—posterior beliefs—is affected by both treatment and prior beliefs. We define mediator as a function denoted by $M(Z;X)$. [^7] Note that potential outcomes are often written as a function of only the manipulated treatment, e.g., $M(Z)$ and $Y(Z)$, to reflect “no causation without manipulation” (Holland, 1986: p. 959). We choose to denote covariates $X$ as arguments to both potential outcomes in order to clarify the structure of covariates and mediators in our potential outcomes. It represents the potential outcomes of causal mediator given treatment $Z$ and covariates $X$. Finally, we denote a potential outcome by $Y(Z,M;X)$.
@@ 3.1 Causal Effects
In order to understand when HTEs can allow for detection of mechanisms, we formalize existing informal conventions. To do so, we decompose the total effect (on a unit, $i$) into direct and indirect effects, as is standard in the causal mediation literature (Imai and Yamamoto, 2013). Suppose that there exist $J\geq 1$ mediators (or mechanisms), indexed by $M_{1},...,M_{J}$. Given two treatment values, $z,z^{\prime}\in Z$, the total effect of $Z$ on $Y$ is:
Our notation varies slightly from conventional presentations of mediation that only consider two mechanisms (mediators) (Imai, Keele, and Tingley, 2010). To this end, $j-$ and $j+$ denote index $h\in J$ such that $h<j$ and $h>j$ respectively. Treatment effects may consist of direct ($DE$) and indirect ($IE_{j}$) effects, as follows: [^8] The intuition for our notation is as follows: define $IE_{0}(z,z^{\prime};X)=DE(z,z^{\prime};X)$. This implies that the first term of $IE_{j}(z,z^{\prime};X)$ and the second term of $IE_{j-1}(z,z^{\prime};X)$ cancel out.  [^9] See Acharya, Blackwell, and Sen (2016) for the identification of the controlled direct effect.
$Y(z^{\prime},M_{j-}(z^{\prime};X),M_{j}(z^{\prime};X),M_{j+}(z;X);X)$
The direct effect, $DE(z,z^{\prime};X)$ represents the direct effect of $Z$ on $Y$ holding mediators at potential outcomes $M_{j}(z;X)$. It is not necessary to believe that treatments produce unmediated (direct) effects on outcomes to use this framework. We allow for direct effects in the interest of generality. The indirect effects, measures the effect on the outcome that operates by changing the potential outcome of the mediator. As is standard, we can then re-write the total effect as follows: [^10] Although there exists multiple ways to decompose total effect and this decomposition holds mechanisms to be independent, all results in the paper hold if other mediators (other than mediator $j$ of interest) are arbitrarily correlated. See additional discussions of correlated mechanisms in the D.
@@
$=E_{X}[DE(z,z^{\prime};X)+\sum_{j=1}^{J}IE_{j}(z,z^{\prime};X)]$ (13)
$=ADE(z,z^{\prime})+\sum_{j=1}^{J}AIE_{j}(z,z^{\prime})$ (14)
@@
HTEs exist with respect to pre-treatment covariate $X_{k}$ if $CATE_{Y}(X_{k}=x)\neq CATE_{Y}(X_{k}=x^{\prime})$ for some $x\neq x^{\prime}\in X_{k}$  [^11] Precisely, the probability measure for the set that contains such $x$ and $x^{\prime}$ is non-zero..
With this framework, it is now possible to express the research question more precisely. First, recall our research question: “under what conditions do HTEs with respect to pre-treatment covariates provide evidence of mechanism activation?” By mechanism activation, we mean that the indirect effect of a mechanism $j$ is non-zero for some unit. The use of heterogeneous treatment effects registers the expectation that the mechanism need not be active for all units in the population. Our question then can be stated more precisely: “Under what conditions are HTEs with respect to a covariate $X_{k}$ sufficient to show that there exists some unit for which $IE_{j}(z,z^{\prime};X_{k})\neq 0$?”
@@ 3.2 Mechanisms and Outcomes
In standard discussions of research design, researchers generally do not make methodological distinctions between different types of outcomes. We introduce a distinction between outcomes that are directly affected and indirectly affected by a mechanism. In the theoretical model of our motivating example and many related theories in this literature, the mechanism—voter learning—acts directly on voter utility. The mechanism indirectly affects vote choice through its (assumed) effect on utility. Definition 3 shows that the distinction between directly- and indirectly-affected outcomes rests on fundamentally on the theorized causal sequence of different outcomes of interest. The sequencing of multiple outcomes has widespread implications for efforts to identify and measure causal effects, but has received limited attention in the existing literature (see Slough, 2022: for an exception).
@@
Given treatment $Z$, let variable $\widetilde{Y}(Z)$ be a (potential) outcome  [^12] We omit M and X in the notation.. If
@@
The characterization of directly- and indirectly-affected outcomes is distinct from whether an outcome is or could be observed by the analyst. As such, our definition is distinct from the concepts of latent and observed outcomes (see, e.g., Fariss, Kenwick, and Reuning, 2020). In principle, directly-affected and indirectly-affected variables could both be latent and/or observed. We illustrate these possibilities concretely in C using three examples from (hypothetical) studies with some type of learning or updating mechanism. One example follows our theoretical motivation. In most studies of exogenous shocks and pro-incumbent voting, researchers measure vote choice but not expected utility. In principle, however, expected utility from a candidate could be measured by eliciting a voter’s willingness to pay for that candidate. In this case, the directly-affected outcome would be observed (non-latent). Vote choice is both indirectly-affected under the model and observed. This example shows that directly- and indirectly-affected outcomes are different from latent and observed outcomes.
The distinction between directly- and indirectly-affected outcomes represents a theoretical commitment about the causal sequencing of outcomes. In contrast, the distinction between latent and observed outcomes represents both a theoretical commitment about the relationship between outcomes (though not necessarily the sequence) and empirical commitments about what outcomes are or could be observed by the researcher. As we will highlight, the distinction directly- and indirectly-affected matters for the use of HTEs to provide evidence of mechanism activation. When outcomes later in a sequence are produced by a (non-zero) linear function of previous outcomes, they can be treated as directly-affected outcomes. [^13] This accomodates popular linear rescalings like $Z$-score transformations of observed outcomes.
@@ 4 HTEs and Mechanisms
When do heterogeneous treatment effects with respect to some covariate, $X_{k}$, provide evidence that a mechanism is active? To answer this question, we must first operationalize the concept of a theoretical mechanism. We view mediators—whether measured or unmeasured—as representations of a theoretical mechanism. For example, in the earlier example about exogenous shocks and voter behavior, the voter’s posterior represents the mediator. If a mediator is measured, researchers can, in theory, use off-the-shelf estimators to estimate the direct and indirect effects described above. Yet, there are two substantial limitations of mediation analysis. First, researchers may not have the ability to measure mediators. Not all mechanisms are measurable and even fewer are measured. Second, mediation analysis typically relies upon an additional ignorability assumption for identification. In our notation, this ignorability assumption, $Y_{i}(z^{\prime},m;X)\perp M_{i}(z;X)\mid Z_{i};X_{i}$, holds that conditional on treatment and pre-treatment covariates $X_{i}$, potential outcomes are independent of the potential outcomes of the mediator (Imai, Keele, and Tingley, 2010). Critics allege that this assumption is unlikely to obtain and show that violation of this assumption generates bias in estimates of indirect effects (Gerber and Green, 2012). [^14] Strezhnev, Kelley, and Simmons (2021) provide an alternative instrumental variables-based identification strategy and sensitivity analysis for assessing mediation effects without invocation of sequential ignorability.
@@
The mediation framework allows us to precisely characterize HTEs with respect to covariates. However, it does not yet provide enough structure to link HTEs (or lack thereof) to mechanisms. To do so, we develop the concept of a mechanism indicator variable (MIV) and impose two assumptions. A MIV for a given mechanism induces a differential (average) causal effects through the mechanism of interest. This can be expressed in terms of (average) indirect effects. To economize notation, we will denote $AIE_{j}(X_{k}=x)=AIE_{j}(z,z^{\prime};X_{k}=x,X_{\neg k})$ as the average indirect effect of mechanism (mediator) $j$ when $X_{k}=x$.
Definition 4 (Mechanism indicator variable (MIV)).
@@
A pre-treatment covariate $X_{k}$ is a mechanism indicator variable for mechanism $j$ if for some $x,x^{\prime}\in X_{k}$, $AIE_{j}(X_{k}=x)\neq AIE_{j}(X_{k}=x^{\prime})$.
@@
We then denote $\textbf{X}_{j}^{MIV}$ as the (possibly empty) set of all possible covariates that satisfy Definition 4 for mechanism $j$. Intuitively, if $X_{k}\in\textbf{X}_{j}^{MIV}$, then covariate $X_{k}$ can serve as an indicator for a mechanism/mediator of interest. [^15] A slightly stronger version of Definition 4 holds when $Y$ is continuously differentiable with respect to $M_{j}$ and $Z$ In this case, Definition 4 can be expressed as $\frac{\partial}{\partial X_{k}}\left(\frac{\partial Y}{\partial M_{j}}\frac{\partial M_{j}}{\partial Z}\right)\neq 0$. Under our definition of MIVs, it could be the case that $X_{k}$ moderates the effect of the treatment on the mediator ($M_{j}$). Interestingly, it could also be the case that $X_{k}$ moderates the effect of the mediator on the outcome. Both possibilities are depicted in Figure 2. The detection of mechanisms using treatment-by-covariate interactions requires researchers to postulate a $MIV$ for a given mechanism.
@@
Figure 2: The two panels depict the causal structure of two MIVs for mechanism $M_{1}$ graphically. Either panel is consistent with Definition 4.
However, postulating a $MIV$ is not sufficient to use HTEs to detect the activation of a mechanism. Specifically, we must also be concerned with whether a given covariate is a $MIV$ for other mediators in addition to $M_{j}$. Specifically, in order to do this, researchers need two additional exclusion restrictions which limit the number of mediators that are affected by a given covariate, $X_{k}$. Logically, if a single covariate moderates multiple mechanisms, we cannot use heterogeneity in that covariate in order to isolate our mechanism of interest ($M_{j}$). Assumption 1 rules out the possibility that direct effects depend on $X_{k}$ and Assumption 2 rules out the possibility that the indirect effects produced by other mechanisms are moderated by $X_{k}$. Importantly, these assumptions do not rule out a direct path from $X_{k}$ to the outcome $Y$. Nor do they rule out a direct path between $X_{k}$ and any other moderator, $M_{\neg j}$, so long as $X_{j}$ does not moderate the effect of treatment through that mechanism.
Assumption 1 (Exclusion I).
@@
Assumption 2 (Exclusion II).
Assumptions 1-2 constrain the relationship between a moderator, $X_{k}$, and other mechanisms/the direct effect. Figure 3 depicts violations of Assumptions 1-2 graphically for a MIV ($X^{MIV}$) of mechanism 1, $M_{1}$. Neither assumption rules out a direct causal relationship between $X_{k}$ and outcome $Y$. In the DAG, such a relationship is depicted in the arrow from $X_{k}$ to $Y$. Assumption 1 rules out the blue dashed path from $X^{MIV}$ to the direct effect of $Z$ on $Y$. Assumption 2 rules out either/both red dot-dashed path from $X^{MIV}$ to the effect of $Z$ on a different mechanism, $M_{2}$. This means that $X_{k}$ does not moderate the effect of $Z$ on $M_{2}$. Nor does $X_{k}$ moderate the effect of $M_{2}$ on $Y$. However, there can be a direct arrow from $X_{k}$ to $M_{2}$. This means that $X_{k}$ can cause (or predict) the level of $M_{2}$, but it cannot interact with treatment in any way to cause $M_{2}$ or the effect of $M_{2}$ on $Y$. These assumptions are crucial for the use of HTEs to detect mechanisms. In their absence, we cannot link a heterogeneous treatment effect to a unique theoretical mechanism representation (mediator) because we cannot ensure additive separability of the effect of the mechanism of interest from the effects of other (possible) mechanisms.
[…]
Figure 3: Assumption 1 rules out the blue dashed path. Assumption 2 rules out both of the red dot-dashed paths. All black solid paths are permissible under Assumptions 1 and 2.
Use of HTEs in order to learn about mechanism activation represents one alternative to mediation analysis. Assumptions 1-2 form the core assumptions underpinning this use of HTE. The comparison to mediation invites a comparison of these exclusion assumptions to the assumption of sequential ignorability. It is useful to note that there is no logical ordering of the two types of assumptions: the exclusion assumptions do not imply sequential ignorability, nor does sequential ignorability imply the exclusion assumptions. This means that HTEs cannot said to be a “more agnostic” or “less agnostic” quantitative test of mechanism activation than mediation. [^16] Mediation analysis attempts to estimate the influence of different mechanisms by decomposing total effects into other causal estimands. To the extent that we care about activation of a mechanism (as in HTEs analysis), we can say that a mechanism is activated in mediation analysis if its indirect effect is non-zero. In some applications, one set of assumptions may be more plausible or defensible than the other, but we cannot make a general claim about the strength of these distinct sets of assumptions. We provide a broader discussion comparing the use of HTEs to mediation analysis in D.

### v2

@@ 3 Framework
We introduce a framework that we use to analyze the relationship between HTEs—as estimated by treatment-covariate interactions—and the detection of causal mechanisms. Our framework is built upon the potential outcomes framework or Neyman-Rubin causal model (Neyman, 1923; Rubin, 1974). We denote treatment by $Z$. In order to consider HTEs with respect to pre-treatment moderators, denote the vector of measured pre-treatment covariates by $X=(X_{1},X_{2},...,X_{K})\in\mathbb{R}^{K}$. Clearly, it need not be the case that all $K$ covariates moderate a given treatment effect.
A valid mediator, or mechanism representation, should: (1) be affected by treatment, $Z$, and (2) have a non-zero effect on the outcome. Importantly, the effect of treatment on a mediator or the mediator’s effect on the outcome could vary with some covariate(s), $X_{k}$. We define a mediator as a function denoted by $M(Z;X)$. [^7] Note that potential outcomes are often written as a function of only the manipulated treatment, e.g., $M(Z)$ and $Y(Z)$, to reflect “no causation without manipulation” (Holland, 1986: p. 959). We choose to denote covariates $X$ as arguments to both potential outcomes in order to clarify the structure of covariates and mediators in our potential outcomes. It represents the potential outcomes of causal mediator given treatment $Z$ and covariates $X$. Finally, we denote a potential outcome by $Y(Z,M;X)$.
@@ 3.1 Causal Effects
In order to understand when HTEs can allow for detection of mechanisms, we formalize existing informal conventions. To do so, we decompose the total effect (on a unit, $i$) into direct and indirect effects, as is standard in the causal mediation literature (Imai and Yamamoto, 2013). Let $M$ be the mediator of interest. We analyze the more general case of multiple mmediators in C. Given two treatment values, $z,z^{\prime}\in Z$, the total effect of $Z$ on $Y$ is:
$TE(z,z^{\prime};X)$ $=Y(z,M(z;X);X)-Y(z^{\prime},M(z^{\prime};X);X)$ (8)
Treatment effects consist of direct ($DE$) and indirect ($IE$) effects, as follows:  [^8] See Acharya, Blackwell, and Sen (2016) for the identification of the controlled direct effect.
$DE(z,z^{\prime};X)=Y(z,M(z;X);X)-Y(z^{\prime},M(z;X);X)$ (9)
$IE(z,z^{\prime};X)=Y(z^{\prime},M(z;X);X)-Y(z^{\prime},M(z^{\prime};X);X)$ (10)
The direct effect, $DE(z,z^{\prime};X)$ represents the direct effect of $Z$ on $Y$ holding mediator $M$ at potential outcomes $M(z;X)$. Our preferred interpretation of the direct effect is the composite effect of all other mechanisms (aside from $M$). However, the direct effect could also include unmediated effects of treatment on an outcome. The indirect effect measures the effect on the outcome that operates by changing the potential outcome of mediator $M$. As is standard, we can then re-write the total effect as follows: [^9] Here, we assume that the direct and indirect effects do not vary at different levels of Z. See more discussion by Imai et al. (2011). All results in the paper hold with other decompositions. See additional discussions of general cases in the C and correlated mechanisms in the E.
$TE(z,z^{\prime};X)$ $=DE(z,z^{\prime};X)+IE(z,z^{\prime};X),$ (11)
@@
$=E_{X}[DE(z,z^{\prime};X)+IE(z,z^{\prime};X)]$ (13)
$:=ADE(z,z^{\prime})+AIE(z,z^{\prime})$ (14)
@@
HTEs exist with respect to pre-treatment covariate $X_{k}$ if $CATE(X_{k}=x)\neq CATE(X_{k}=x^{\prime})$ for some $x\neq x^{\prime}\in X_{k}$. [^10] More precisely, the probability measure for the set that contains such $x$ and $x^{\prime}$ is non-zero.
With this framework, it is now possible to express the research question more precisely. First, recall our research question: “under what conditions do HTEs with respect to pre-treatment covariates provide evidence of mechanism activation?” By mechanism activation, we mean that the indirect effect of a mechanism is non-zero for some unit. The use of heterogeneous treatment effects registers the expectation that the mechanism need not be active for all units in the population. Our question then can be stated more precisely: “Under what conditions are HTEs with respect to a covariate $X_{k}$ sufficient to show that there exists some unit for which $IE(z,z^{\prime};X_{k})\neq 0$?”
@@ 3.2 Mechanisms and Outcomes
In standard discussions of research design, researchers generally do not make methodological distinctions between different types of outcomes. We introduce a distinction between outcomes that are directly affected and indirectly affected by a mechanism. In the theoretical model of our motivating example and many related theories in this literature, the mechanism—voter learning—acts directly on voter utility. The mechanism indirectly affects vote choice through its (assumed) effect on utility. Definition 3 shows that the distinction between directly- and indirectly-affected outcomes rests on fundamentally on the theorized causal sequence of different outcomes of interest. The sequencing of multiple outcomes has widespread implications for efforts to identify and measure causal effects, but has received limited attention in the existing literature (for one exception, see Slough, 2023).
@@
Given treatment $Z$, let variable $\widetilde{Y}(Z)$ be a (potential) outcome. If
@@
The characterization of directly- and indirectly-affected outcomes is distinct from whether an outcome is or could be observed by the analyst. As such, our definition is distinct from the concepts of latent and observed outcomes (e.g., Fariss, Kenwick, and Reuning, 2020). In principle, directly-affected and indirectly-affected variables could both be latent and/or observed. We illustrate these possibilities concretely in D using three examples from (hypothetical) studies with some type of learning or updating mechanism.
The distinction between directly- and indirectly-affected outcomes represents a theoretical commitment about the causal sequencing of outcomes. In contrast, the distinction between latent and observed outcomes represents both a theoretical commitment about the relationship between outcomes (though not necessarily the sequence) and empirical concerns about what outcomes are or could be observed by the researcher. As we will highlight, the distinction directly- and indirectly-affected matters for the use of HTEs to provide evidence of mechanism activation. When outcomes later in a sequence are produced by a (non-zero) linear function of previous outcomes, they can be treated as directly-affected outcomes. This accomodates popular linear rescalings like $Z$-score transformations of observed outcomes.
@@ 4 HTEs and Mechanisms
When do heterogeneous treatment effects with respect to some covariate, $X_{k}$, provide evidence that a mechanism is active? To answer this question, we must first operationalize the concept of a theoretical mechanism. We view mediators—whether measured or unmeasured—as representations of a theoretical mechanism. For example, in the earlier example about exogenous shocks and voter behavior, the voter’s posterior belief represents the mediator. If a mediator is measured, researchers can, in theory, use off-the-shelf estimators to estimate the direct and indirect effects described above. Yet, there are two substantial limitations of mediation analysis. First, researchers may not have the ability to measure mediators. Not all mechanisms are measurable and even fewer are measured. Second, mediation analysis typically relies upon an additional ignorability assumption for identification. In our notation, this ignorability assumption, $Y_{i}(z^{\prime},m;X_{i})\perp M_{i}(z;X_{i})\mid Z_{i}=z;X_{i}$, holds that conditional on treatment and pre-treatment covariates $X_{i}$, potential outcomes are independent of the potential outcomes of the mediator (Imai, Keele, and Tingley, 2010). Critics allege that this assumption is unlikely to obtain and show that violation of this assumption generates bias in estimates of indirect effects (Gerber and Green, 2012). [^11] Strezhnev, Kelley, and Simmons (2021) relax sequential ignorability through an instrumental variables-based identification strategy and sensitivity analysis for assessing mediation effects.
@@
The mediation framework allows us to precisely characterize HTEs with respect to covariates. However, it does not yet provide enough structure to link HTEs (or lack thereof) to mechanisms. To do so, we develop the concept of a mechanism detector variable (MDV) and impose an assumption. A MDV for a given mechanism induces a differential (average) causal effects through the mechanism of interest on the directly-affected outcome. This can be expressed in terms of (average) indirect effects. To economize notation, we will denote $AIE(X_{k}=x)=AIE(z,z^{\prime};X_{k}=x,X_{\neg k})$ as the average indirect effect of mechanism (mediator) when $X_{k}=x$.
Definition 4 (Mechanism detector variable (MDV)).
@@
A pre-treatment covariate $X_{k}$ is a mechanism detector variable for mechanism $M$ if for some $x,x^{\prime}\in X_{k}$, $AIE(X_{k}=x)\neq AIE(X_{k}=x^{\prime})$ for the directly-affected outcome.
@@
We then denote $\textbf{X}^{MDV}$ as the (possibly empty) set of all possible covariates that satisfy Definition 4 for the mechanism $M$ of a directly-affected outcome. Intuitively, if $X_{k}\in\textbf{X}^{MDV}$, then covariate $X_{k}$ can serve as an indicator for a mechanism/mediator of interest. [^12] A slightly stronger version of Definition 4 holds when $Y$ is continuously differentiable with respect to $M$ and $Z$ In this case, Definition 4 can be expressed as $\frac{\partial}{\partial X_{k}}\left(\frac{\partial Y}{\partial M}\frac{\partial M}{\partial Z}\right)\neq 0$. Under our definition of MDVs, it could be the case that $X_{k}$ moderates the effect of the treatment on the mediator. Interestingly, it could also be the case that $X_{k}$ moderates the effect of the mediator on the outcome. Both possibilities are depicted in Figure 2. The detection of mechanisms using treatment-by-covariate interactions requires researchers to postulate a $MDV$ for a given mechanism.
@@
Figure 2: The two panels depict the causal structure of two MDVs for mechanism $M$ graphically. Both panels are consistent with Definition 4.
However, postulating a $MDV$ is not sufficient to use HTEs to detect the activation of a mechanism. Specifically, we must also be concerned with whether a given covariate is a $MDV$ for other mechanisms (incorporated in the direct effect). Specifically, in order to do this, researchers need an exclusion restriction to limit the number of mediators that are affected by a given covariate, $X_{k}$. Logically, if a single covariate moderates a mechanism of interest and the direct effect, we cannot use HTEs in that covariate to isolate the mechanism of interest. Assumption 2 rules out the possibility that direct effects depend on $X_{k}$. In the more general multiple-mechanism case, we also need exclusion assumptions to rule out the possibility that the indirect effects produced by other mechanisms are moderated by $X_{k}$ (see C).
Assumption 1 (Exclusion).
@@
Assumption 2 constrains the relationship between a moderator, $X_{k}$, and the direct effect (taken to be a composite of other possible mechanims). Figure 3 depicts a violation of Assumption 2 graphically for a MDV ($X_{k}$) of the mechanism $M$. Assumption 2 rules out the blue dashed path from $X_{k}$ to the direct effect of $Z$ on $Y$. Given our notation of interactions in a DAG, this means that $X_{k}$ does not moderate the direct effect of $Z$ on $Y$. This assumption is necessary for the use of HTEs to detect mechanisms. In its absence, we cannot link a heterogeneous treatment effects to a unique theoretical mechanism representation (here, $M$) because we have no way to ensure additive separability of the effect of $M$ from the effects of other mechanisms. Importantly, Assumption 2 does not imply that $X_{k}$ must be independent of the direct effect of $Z$ on $Y$; these quantities can be arbitraily correlated so long as $X_{k}$ does not induce heterogeneity in the direct effect.
$Z$$M$$Y$$X_{k}$
Figure 3: Assumption 2 rules out the blue dashed path. All black and orange paths are permissible under Assumption 2. For $X_{k}$ to be a $MDV$, at least one of the orange paths must be present.
Use of HTEs in order to learn about mechanism activation represents one alternative to mediation analysis. Assumptions 2 forms the core assumption underpinning this use of HTE. The comparison to mediation analysis invites a comparison of the exclusion assumption to the assumption of sequential ignorability. It is useful to note that there is no logical ordering of the two types of assumptions: the exclusion assumption does not imply sequential ignorability, nor does sequential ignorability imply the exclusion assumption. This means that HTEs cannot said to be a “more agnostic” or “less agnostic” quantitative test of mechanism activation than mediation. [^13] Mediation analysis attempts to estimate the influence of different mechanisms by decomposing total effects into other causal estimands. When our primary interest is the activation of a mechanism (as in HTEs analysis), we can say that a mechanism is activated in mediation analysis if its indirect effect is non-zero. In some applications, one set of assumptions may be more plausible or defensible than the other, but we cannot make a general claim about the strength of these distinct sets of assumptions. We provide a broader discussion comparing the use of HTEs to mediation analysis in E.